OpenStudy (rosales0228):

GIVING MEDAL!!!!!!! Which expression is equivalent to sin2x−sinx−2sinx−2 ?

OpenStudy (rosales0228):

sin x + 1 sin x − 1 sin2x -sin2x

OpenStudy (solomonzelman):

is the first thing here `sin²x` or `sin(2x)` ?

OpenStudy (solomonzelman):

when you wrote sin2x is it sin²x or sin(2x) ? Hello ?

OpenStudy (oleg3321):

solomon plrase help me next

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \sin^2x−\sin(x)−2\sin(x)− 2 }\) \(\normalsize\color{blue}{ \sin x(\sin(x)-1)−2(\sin(x)− 1) }\) \(\normalsize\color{blue}{ (\sin x−2)(\sin x− 1) }\)

OpenStudy (solomonzelman):

I am in trouble... are you sure you wrote the question correctly ?

OpenStudy (rosales0228):

yeah I copied and pasted it

OpenStudy (solomonzelman):

you messaged me, saying `sin²(-x)` so it is not \(\normalsize\color{blue}{\sin 2(-x)−\sin x−2\sin x−2 }\) , right?

OpenStudy (solomonzelman):

Just \(\normalsize\color{blue}{\sin^2x−\sin x−2\sin x−2 }\) ?